OpenStudy (anonymous):

How do you find the slope of y=4?

OpenStudy (anonymous):

it's a horizontal line, since there is no function of x. Therefore it's slope is zero!

Directrix (directrix):

You could use the slope formula to "see" this algebraically. Two points on the line y = 4 could be ( 1, 4 ) and ( 6 , 4). The only requirement is that the y-coordinate be 4. slope = (4 - 4) / (1-6) = 0/(-5) = 0 which can be thought of as 0/1. Each time x increases by 1 unit, y increases by 0 units. Hence, a horizontal line.